Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-price
Answer #1

a.

Average assets                                       3,000,000.00
Turnover Ratio 0.6
Formula Net Sales/Average total assets
ie., 0.6 = Net sales/3000000
So Net Sales                                       1,800,000.00
Gross Margin 40% (given)
So,                                           720,000.00
Hence, Current Sales                                       1,800,000.00
ROI = 720,000/3000000*100 24%

b.Since ROI and Average Total Assets are the same, Margin will be = 3000,000*24% = 720,000. This is 30% of sales. So Sales => 720,000/30*100 = 2400000.

c. With reduced prices total sales has to be increased from $ 1800,000 to $ 2400000. Ie increase of $ 600,000. or 600,000/1800,000*100 = 33% and not 50%.

d. Marketing strategy involves costs usually fixed in nature viz salaries, promotion, advertisement costs or variable like sales commission and volume based incentives. Hence, such marketing should boost up the contribution to that extent It recovers the variable costs then the fixed marketing costs and provide for 40% on sales margin and finally maintain ROI 30% on assets.

e. Charlie can focus on cutting down his costs. Presently 60% of sales are his costs (100% - 40% margin). Cutting down costs can also help him in reducing prices thus generate higher volume of sales. The margins due to reduced costs also can be maintained easily even though prices are reduced.